# R-Judge: Benchmarking Safety Risk Awareness for LLM Agents

### v2

@@ D.2 Model Failures
D.2.1 GPT-4 Failures
Here we illustrate 3 key capability flaws of GPT-4 discussed in Section 5.2, including flaws of scenario simulation (Figure 6), understanding adaptability (Figure 7), and safety alignment (Figure 8). Distributions are shown in Figure 4.
Figure 4: Proportion of 3 key GPT-4 capability flaws with Zero-Shot-CoT in safety judgment test.
(i) Scenario Simulation: Fail to retrieve relevant knowledge and reason in specific scenarios. Some safety risks are hidden in the complex multi-turn agent interaction, the identification of which demands LLM monitors to infer the effect of certain actions, such as the effect of link sharing with edit access. Due to the lack of scenario knowledge or reasoning ability, GPT-4 may fail to associate corresponding knowledge and reason in specific scenarios to identify risks. For example, GPT-4 failed in the case where the agent shared the file link on Twitter with edit access instead of comment access.
(ii) Understanding Adaptability: Unable to comprehend risks in specific conditions. Risks are closely tied to specific conditions, i.e., risky actions in Scenario A may be safe in Scenario B instead. Despite possessing a comprehensive understanding of security concerns through safety alignment, GPT-4 can sometimes exhibit rigidity and excessive concern due to its limited comprehension of certain conditions. For instance, in the case of false positives, GPT-4 may respond by stating that "the actions taken by the agent are generally safe, but there are a few potential security concerns…" and incorrectly categorize them as unsafe. Conversely, in the case of false negatives, GPT-4 may recognize that the agent should seek confirmation from the user regarding a potentially risky action but mistakenly classify it as safe.
Models Safety Judgment
F1 Recall Specificity Validity
Human 89.07 82.17 95.76 100.00
100.00 100.00 100.00 100.00
GPT-4 72.52 62.00 83.64 100.00
99.50 100.00 98.18 100.00
Llama-2-13b-chat-hf 38.86 34.00 25.45 50.97
96.00 96.00 92.73 100.00
Vicuna-13b-v1.5-16k 43.24 32.00 70.91 99.35
93.07 94.00 85.45 100.00
Vicuna-7b-v1.5-16k 36.88 26.00 72.73 100.00
92.78 90.00 92.73 100.00
ChatGPT 39.42 27.00 81.82 100.00
91.87 96.00 76.36 100.00
Vicuna-7b-v1.5 19.35 12.00 78.18 100.00
81.66 69.00 100.00 100.00
Vicuna-13b-v1.5 30.30 20.00 78.18 100.00
68.42 65.00 54.55 65.16
Random 50.00 50.00 50.00 50.00
Llama-2-7b-chat-hf 21.56 18.00 10.91 37.42
24.84 20.00 25.45 36.13
Table 7: Main Results(%) of the Oracle Test (Safety Judgment w/ Risk Description). In each segment, the first line lists baseline results of Safety Judgment (w/o Risk Description), and the second line lists results of Oracle Test (Safety Judgment w/ Risk Description). Reference Scores, including Random as baseline and Human as upper bound, are marked as underline.
D.2.2 Llama-2 Failures
In the safety judgment test, the Llama-2 series present instability to output ’safe’ or ’unsafe’. We show a representative failure of Llama-2-13b-chat-hf in Figure 5.
Appendix E Experiment Resources & Cost
In this section, we will report the computational budget and computing infrastructure utilized in our experiments to facilitate reproduction. The time required for the Zero-Shot-CoT and Few-Shot-CoT experiments is similar, as both require two rounds of inference for each sample. The Oracle Test requires only one round and thus needs almost the half time of them. We report statistics of the main Zero-Shot-CoT experiment for reference.
In the case of API-based models, the time required per experiment (traversing all samples in one pass) in R-Judge is generally under 1 hour, although it may be influenced by the state of the network.
As for the open-sourced models, we employed a single A100 GPU for inference. The time consumption for each model in each experiment is approximately 1.5 hours.

### v3

@@ E.4 Model Failures
E.4.1 GPT-4o Failures
Here we illustrate 3 key capability flaws of GPT-4o discussed in Section 5.3, including flaws of scenario simulation (Figure 11), understanding adaptability (Figure 12), and safety alignment (Figure 13).
E.4.2 Llama-2 Failures
In the safety judgment test, the Llama-2 series presents instability to output ‘safe’ or ‘unsafe’. We show a representative failure of Llama-2-13b-chat-hf in Figure 10.
Appendix F Experiment Settings
F.1 Hyperparameter
For paramaters including temperature and top_p, we use the default values of official model releases. And we use FastChat to deploy models with half-precision weights.
F.2 Experiment Resources & Cost
In this section, we report the computational budget and computing infrastructure utilized in our experiments to facilitate reproduction. The time required for the Zero-Shot-CoT and Few-Shot-CoT experiments is similar, as both require two rounds of inference for each sample. We report statistics of the main Zero-Shot-CoT experiment for reference.
In the case of API-based models, the time required per experiment (traversing all 569 samples in one pass) in R-Judge is generally under 1 hour, although it may be influenced by the state of the network.
As for the open-sourced models, we employed a single A100 GPU for inference. The time consumption for each model in each experiment is approximately 1.5 hour.
F.3 Model Information
Table 9 lists concrete information about models in the experiments.
Model Model Size Access Version Creator
GPT-4o undisclosed api gpt-4o-2024-05-13 OpenAI
ChatGPT undisclosed api gpt-3.5-turbo-0125
Llama-2-13b-chat-hf 13B weights - Meta
Llama-2-7b-chat-hf 7B weights -
LlamaGuard-7b 7B weights -
Meta-Llama-3-8B-Instruct 8B weights -
Meta-Llama-Guard-2-8B 8B weights -
Vicuna-13b-v1.5 13B weights v1.5 LMSYS
Vicuna-13b-v1.5-16k 13B weights v1.5
Vicuna-7b-v1.5 7B weights v1.5
Vicuna-7b-v1.5-16k 7B weights v1.5
Mistral-7B-Instruct-v0.2 7B weights v0.2 Mistral AI
Mistral-7B-Instruct-v0.3 7B weights v0.3
Table 9: LLMs evaluated in this paper.